OpenStudy (anonymous):

If you were solving a system of equations and you came to a statement like 4 = 4, what do you know about the system?

OpenStudy (amistre64):

when does 4=4?

OpenStudy (anonymous):

Umm, I'm not sure..

OpenStudy (amistre64):

4=4 is a true statement and is ALWAYS true. this should bring to mind that the system ALWAYS shares common points, so they are congruent systems.

OpenStudy (amistre64):

congruent is prolly not technically correct ... but it means they are the same system regardless.

OpenStudy (anonymous):

I have an equation, can you see if I got the right answer real quick?

OpenStudy (amistre64):

prolly

OpenStudy (anonymous):

Solve the following system and show all of your work. 3y − 2x = 11 y + 2x = 9 I got : x=2, y=5

OpenStudy (amistre64):

you didnt show your work ... but as a dbl chk: 3y - 2x = 11 5 2 ----------- 15 - 4 = 11 is true y + 2x = 9 5 2 ---------- 5 + 4 = 9 is true

OpenStudy (anonymous):

Thank you so much!

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

The graph of a system of equations with the same slope and the same y-intercepts will have no solutions. Always Sometimes Never @amistre64